# Latent Factor Analysis in Short Panels

### v2

@@ Abstract
We develop inferential tools for latent factor analysis in short panels. The pseudo maximum likelihood setting under a large cross-sectional dimension $n$ and a fixed time series dimension $T$ relies on a diagonal $T\times T$ covariance matrix of the errors without imposing sphericity nor Gaussianity. We outline the asymptotic distributions of the latent factor and error covariance estimates as well as of an asymptotically uniformly most powerful invariant (AUMPI) test for the number of factors based on the likelihood ratio statistic. We derive the AUMPI characterization from inequalities ensuring the monotone likelihood ratio property for positive definite quadratic forms in normal variables. An empirical application to a large panel of monthly U.S. stock returns separates month after month systematic and idiosyncratic risks in short subperiods of bear vs. bull market based on the selected number of factors. We observe an uptrend in the paths of total and idiosyncratic volatilities while the systematic risk explains a large part of the cross-sectional total variance in bear markets but is not driven by a single factor. Rank tests show that observed factors struggle spanning latent factors with a discrepancy between the dimensions of the two factor spaces decreasing over time.
Keywords: Latent factor analysis, uniformly most powerful invariant test, panel data, large $n$ and fixed $T$ asymptotics, equity returns.
JEL codes:
C12, C23, C38, C58, G12.
@@ 1 Introduction
Latent variable models have been used for a long time in econometrics (Aigner et al. (1984)). Here, we study large cross-sectional latent factor models with small time dimension. Two common methods for estimation of latent factor spaces are principal component analysis (PCA) and factor analysis (FA), see Anderson (2003) Chapters 11 and 14. They cover multiple applications in finance and economics as well as in social sciences in general. They are often used in exploratory analysis of data. Omitted latent factors are also called interactive fixed effects in the panel literature (Pesaran (2006), Bai (2009), Moon and Weidner (2015), Freyberger (2018)). In recent work, Fortin, Gagliardini and Scaillet (FGS, 2023) show how we can use PCA to conduct inference on the number of factors in such models without making Gaussian assumptions. Their methodology relies on sphericity of the idiosyncratic variances since this restriction is both necessary and sufficient for consistency of latent factor estimates with small $T$ (Theorem 4 of Bai (2003)). FGS provides a discussion of the (in)consistency of the PCA estimator with fixed T from the vantage point of the well-known incidental parameter problem of the panel data literature (Neyman and Scott(1948); see Lancaster (2000) for a review). In PCA, sphericity allows to identify the number $k$ of factors from the $k$ first eigenvalue spacings being larger than zero, and being zero the subsequent ones. On the contrary, the FA strategy does not exploit eigenvalue spacings and does not require sphericity. However, inference with small $T$ up to now mostly relies on (often restrictive) assumptions such as Gaussian variables (with a notable exception by Anderson and Amemiya (1988)) and error homoskedasticity across sample units. Those are untenable assumptions in our application with stock returns. The strong assumption of sphericity might also fail to hold in some samples. If it happens, our Monte Carlo experiments with non-Gaussian errors show that the eigenvalue spacing test of FGS exhibit size distortions of over 80 percentage points for a nominal size of 5%. [^1] That massive over-rejection translates into an average estimated number of factors often above 10 instead of 2 (true value) in our simulation design when $T=24$ (a sample size close to $T=20$ in our empirics). We get similar Monte Carlo results for a constrained Likelihood Ratio test when sphericity does not hold (see Section 4.3). Our Monte Carlo experiments also reveal that the classical chi-square test of the FA theory obtained under cross-sectionally homoskedastic Gaussian errors (Anderson (1963)) suffer from massive over-rejection by around 80 percentage points for a nominal size of 5% when $T=24$.
@@
A central and practical issue in applied work with latent factors is to determine the number of factors. For models with unobservable (latent) factors only, Connor and Korajczyk (1993) are the first to develop a test for the number of factors for large balanced panels of individual stock returns in time-invariant models under covariance stationarity and homoskedasticity. Unobservable factors are estimated by the method of asymptotic principal components developed by Connor and Korajczyk (1986) (see also Stock and Watson (2002)). For heteroskedastic settings, the recent literature on large balanced panels with static factors has extended the toolkit available to researchers. A first strand of that literature focuses on consistent estimation procedures for the number of factors. Bai and Ng (2002) introduce a penalized least-squares strategy to estimate the number of factors, at least one. Ando and Bai (2015) extend that approach when explanatory variables are present in the linear specification (see Bai (2009) for homogeneous regression coefficients). Onatski (2010) looks at the behavior of differences in adjacent eigenvalues to determine the number of factors when $n$ and $T$ are both large and comparable. Ahn and Horenstein (2013) opt for a similar strategy based on eigenvalue ratios. Caner and Han (2014) propose an estimator with a group bridge penalization to determine the number of unobservable factors. Based on the framework of Gagliardini, Ossola and Scaillet (2016), Gagliardini, Ossola and Scaillet (2019) build a simple diagnostic criterion for approximate factor structure in large panel datasets. Given observable factors, the criterion checks whether the errors are weakly cross-sectionally correlated, or share one or more unobservable common factors (interactive effects), and selects their number; see Gagliardini, Ossola and Scaillet (2020) for a survey of estimation of large dimensional conditional factor models in finance. A second strand of that literature develops inference procedures for hypotheses on the number of latent factors. Onatski (2009) deploys a characterization of the largest eigenvalues of a Wishart-distributed covariance matrix with large dimensions in terms of the Tracy-Widom Law. To get a Wishart distribution, Onatski (2009) assumes either Gaussian errors, or $T$ much larger than $n$. Kapetanios (2010) uses subsampling to estimate the limit distribution of the adjacent eigenvalues.
@@
This paper puts forward methodological and empirical contributions that complement the above literature. (i) On the methodological side, we extend the inferential tools of FA to non-Gaussian and non-i.i.d. settings. First, we characterize the asymptotic distribution of FA estimators obtained under a pseudo maximum likelihood approach where the time-series dimension is held fixed while the cross-sectional dimension diverges. Hence, the asymptotic analysis targets short panels, and allows for cross-sectionally heteroskedastic and weakly dependent errors. Cochrane (2005, p. 226) argues in favour of the development of appropriate large-$n$ small-$T$ tools for evaluating asset pricing models, a problem only partially addressed in finance. In a short panel setting, Zaffaroni (2019) considers inference for latent factors in conditional linear asset pricing models under sphericity based on PCA, including estimation of the number of factors. [^2] Raponi, Robotti and Zaffaroni (2020) develop tests of beta-pricing models and a two-pass methodology to estimate the ex-post risk premia (Shanken (1992)) associated to observable factors (see Kleibergen and Zhan (2023) for robust-identification inference based a continuous updating generalized method of moments). Kim and Skoulakis (2018) deals with the error-in-variable problem of the two-pass methodology with small $T$ by regression-calibration under sphericity and a block-dependence structure. The small $T$ setting mitigates concerns for panel unbalancedness (outside the straightforward missing-at-random mechanism) and corresponds to a locally time-invariant factor structure accommodating globally time-dependent features of general forms. It is also appealing to macroeconomic data observed quarterly. For the sake of space, we put part of the theory, namely inference for FA estimates, in the Online Appendix (OA). We refer to Bai and Li (2016) for inference when $n$ and $T$ are both large (see Bai and Li (2012) for the cross-sectional independent case). Second, we build on our new theoretical results for FA to develop testing procedures for the number of latent factors in a short panel which rely on neither sphericity nor Gaussianity nor cross-sectional independence, thereby extending tests based on eigenvalues, as in Onatski (2009), to small $T$, and as in FGS, to non-spherical errors, thanks to an FA device. Here, we deliver asymptotic feasible distributional theory even if FA residuals are not consistent estimates of the true errors under fixed $T$. We further derive the Asymptotically Uniformly Most Powerful Invariant (AUMPI) property of the FA likelihood ratio (LR) test statistic in the non-Gaussian case under inequality restrictions on the DGP parameters, and cover inference with weak factors. The AUMPI property is rare and sought-after in testing procedures (see Engle (1984) for a discussion and Romano, Shaikh and Wolf (2010) for a survey of optimality approaches in testing problems), and often holds only under restrictive assumptions such as Gaussianity. We show that the AUMPI property can hold even if the asymptotic distribution of the LR statistic is driven by a weighted average of independent chi-square variates in our context instead of a single chi-square variate. We achieve that by providing novel sufficient conditions so that the ratio of the density under local alternative hypotheses and the null hypothesis satisfies the Monotone Likelihood Ratio (MLR) property. Hence, the FA theory gathered in the below consists in a body of new results including their proofs, and differs completely from PCA theory. (ii) On the empirical side, we apply our FA methodology to panels of monthly U.S. stock returns with large cross-sectional and small time-series dimensions, and investigate how the number of driving factors changes over time and particular periods. Furthermore, month after month, we provide a novel separation [^3] Such a decomposition with PCA estimates is invalid without sphericity because of inconsistency of $\hat{F}$. of the risk coming from the systematic part and the risk coming from the idiosyncratic part of returns in short subperiods of bear vs. bull market based on the selected number of factors. We observe an uptrend in the estimated paths of total and idiosyncratic volatilities (see also Campbell et al. (2023)) while the systematic risk explains a large part of the cross-sectional total variance in bear markets but is not driven by a single latent factor. We also investigate whether standard observed factors span the estimated latent factors using rank tests in order to suit our fixed $T$ setting. Observed factors struggle spanning latent factors with a discrepancy between the dimensions of the two factor spaces decreasing over time.
@@
The outline of the paper is as follows. In Section 2, we consider a linear latent factor model and introduce test statistics on the number of latent factors based on FA. Section 3 presents a feasible asymptotic distributional theory for inference in short panels under a block-dependence structure to allow for weak dependence in the cross-section. Section 4 discusses three special cases, i.e., Gaussian errors (yielding the classical chi-square test of FA theory in the cross-sectionally homoskedastic case), settings where the asymptotic distribution under Gaussian errors still holds for the test statistics, and spherical errors. Section 5 is dedicated to local asymptotic power and AUMPI tests. We provide our empirical application in Section 6 and our concluding remarks in Section 7. Appendices A and B gather the regularity assumptions and proofs of the main theoretical results. Appendix C gives a Monte Carlo assessment of size and power and selection procedure for the number of factors for the LR test. We place all omitted proofs and additional analyses in Appendices D-F in Online Appendix (OA). Appendix G collects the maximum value of $k$ as a function of $T$. Besides, we gather all explicit formulas not listed in the core text but useful for coding in an online “Supplementary Materials for Coding” (SMC) attached to the replication files. We also put there other numerical checks and additional Monte Carlo results to assess the impact of non-sphericity for the eigenvalue spacing test of FGS and the constrained LR test, as well as of using the classical chi-square test on the size and selection procedure for the number of factors.
@@ 2 Latent factor model and likelihood ratio test
where $y_{i}=(y_{i,1},...,y_{i,T})^{\prime}$ and $\varepsilon_{i}=(\varepsilon_{i,1},...,\varepsilon_{i,T})^{\prime}$ are $T$-dimensional vectors of observed data and unobserved error terms for individual $i$. The $k$-dimensional vectors $\beta_{i}=(\beta_{i,1},...,\beta_{i,k})^{\prime}$ are latent individual effects, while $\mu$ and $F$ are a $T\times 1$ vector and a $T\times k$ matrix of unknown parameters. The number of latent factors $k$ is an unknown integer smaller than $T$. In matrix notation, model (1) reads $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+\varepsilon$, where $Y$ and $\varepsilon$ are $T\times n$ matrices, $\beta$ is the $n\times k$ matrix with rows $\beta_{i}^{\prime}$, and $1_{n}$ is a $n$-dimensional vector of ones.
@@
Matrix $V_{\varepsilon}$ is the limit cross-sectional average of the - possibly heterogeneous - errors’ unconditional variance-covariance matrices. The diagonality condition in Assumption 1 is standard in FA (in the more restrictive formulation involving i.i.d. data).
@@
In our empirics with a large cross-sectional panel of returns for $n$ assets over a short time span with $T$ periods, vectors $y_{i}$ and $\varepsilon_{i}$ stack the monthly returns and the idiosyncratic errors of stock $i$. Any row vector $f_{t}^{\prime}:=(f_{t,1},...,f_{t,k})$ of matrix $F$ yields the latent factor values in a given month $t$, and vector $\beta_{i}$ collects the factor loadings of stock $i$. In this finance application, we assume the No-Arbitrage (NA) principle to hold, so that the entries $\mu_{t}$ in the intercept vector in Equation (1) account for the (possibly time-varying) risk-free rate and (possibly non-zero) cross-sectional mean of stock betas.444 Under NA, the intercept term in the asset return model $y_{i}=\mu_{i}+\tilde{F}\tilde{\beta}_{i}+\varepsilon_{i}$ is $\mu_{i}=r_{f}+1_{T}\nu^{\prime}\tilde{\beta}_{i}$, where $r_{f}$ is the $T$-dimensional vector whose entries collect the (possibly time-varying) risk-free rates, $\nu=(\nu_{1},...,\nu_{k})^{\prime}$ is a $k$-dimensional vector of parameters, and $1_{T}$ is a $T$-dimensional vector of ones (see e.g.  Gagliardini, Ossola and Scaillet (2016)). We can absorb term $1_{T}\nu^{\prime}\tilde{\beta}_{i}$ into the systematic part to get $y_{i}=r_{f}+F\tilde{\beta}_{i}+\varepsilon_{i}$ with $F=\tilde{F}+1_{T}\nu^{\prime}$. It holds irrespective of the latent factors being tradable or not. If the factors are tradable, we further have $\nu=0$ from the NA restriction. Akin to standard formulation of FA, we recenter the latent effects by subtracting their mean $\tilde{\mu}_{\tilde{\beta}}=\frac{1}{n}\sum_{i=1}^{n}\tilde{\beta}_{i}$, to get model (1) with $\beta_{i}=\tilde{\beta}_{i}-\tilde{\mu}_{\tilde{\beta}}$ and $\mu=r_{f}+F\tilde{\mu}_{\tilde{\beta}}$. Thus, the linear FA model (1) yields $y_{i,t}=\mu_{t}+f_{t}^{\prime}\beta_{i}+\varepsilon_{i,t}$, that is the standard formulation in asset pricing. We cover the Capital Asset Pricing Model (CAPM) when the single latent factor is the excess return of the market portfolio. Assumption 1 allows for serial dependence in idiosyncratic errors in the form of martingale difference sequences, like individual GARCH and Stochastic Volatility (SV) processes, as well as weak cross-sectional dependence (see Assumption 2 below). It also accommodates common time-varying components in idiosyncratic volatilities by allowing different entries along the diagonal of $V_{\varepsilon}$; see Renault, Van Der Heijden and Werker (2023) for arbitrage pricing in such settings. [^5] When there is a common random component in idiosyncratic volatilities, we have $V_{\varepsilon}=\underset{n\rightarrow\infty}{plim}\frac{1}{n}\varepsilon\varepsilon^{\prime}$ by a suitable version of the Law of Large Number (LLN) conditional on the sigma-field generated by this common component. With fixed $T$, we treat the sample realizations of the common component in idiosyncratic volatilities as unknown time fixed effects (the diagonal elements of matrix $V_{\varepsilon}$), which yields time heterogeneous distributions for the errors. It is how the unconditional expectation in Assumption 1 has to be understood.

### v3

@@ Abstract
We develop a pseudo maximum likelihood method for latent factor analysis in short panels without imposing sphericity nor Gaussianity. We derive an asymptotically uniformly most powerful invariant test for the number of factors. On a large panel of monthly U.S. stock returns, we separate month after month systematic and idiosyncratic risks in short subperiods of bear vs. bull market. We observe an uptrend in the paths of total and idiosyncratic volatilities. The systematic risk explains a large part of the cross-sectional total variance in bear markets but is not driven by a single factor and not spanned by observed factors.
Keywords: Latent factor analysis, uniformly most powerful invariant test, panel data, large $n$ and fixed $T$ asymptotics, equity returns. JEL codes: C12, C23, C38, C58, G12.
@@ 1 Introduction
Latent variable models have been used for decades in econometrics (Aigner et al. (1984)). In this paper, we study large cross-sectional latent factor models for panels of equity returns with small time dimension. Latent factors capture unobserved sources of dependence across the sample units and are largely adopted in multiple applications in finance and economics as well as in social sciences. In asset pricing, they underly the celebrated Arbitrage Pricing Theory (APT) of Ross (1976); see also Chamberlain and Rothschild (1983), Giglio and Xiu (2021), and Zaffaroni (2025). Omitted latent factors are also called interactive fixed effects (Pesaran (2006), Bai (2009), Moon and Weidner (2015), Gobillon and Magnac (2016), Freyberger (2018)) or time-varying individual effects (Ahn et al. (2001)) in the panel literature. We find them in asset embeddings, i.e., numerical representations of asset or firm characteristics in a vector space (Gabaix et al. (2023)).
@@
Two common methods, sometimes put under the same terminology of “factor analysis", for estimation of latent factor spaces are principal component analysis (PCA) and factor analysis (FA), see Anderson (2003) Chapters 11 and 14, and Magnus and Neudecker (2007) Chapter 17, for a clear distinction between the two approaches. PCA relies on minimizing a least squares objective function (leading to a convenient characterization in terms of an eigenvalue-eigenvector decomposition of a variance-covariance matrix) while FA relies on maximizing a Gaussian (pseudo) likelihood function (related to minimizing a weighted least squares objective function as in a weighted PCA) to estimate the latent factors. Notably, FA boils down to PCA under the sphericity condition, i.e., when the $T\times T$ covariance matrix of error terms is a multiple of the identity matrix. Otherwise, PCA and FA estimates generally differ - even asymptotically, when one panel dimension (in our case, $T$) is fixed. In fact, error sphericity is both necessary and sufficient for consistency of PCA latent factor estimates with small $T$ (Theorem 4 of Bai (2003)) and we cannot get rid of this restrictive assumption if we choose to apply PCA to short panels. In recent work, Fortin, Gagliardini and Scaillet (FGS, 2023) show how we can use PCA to conduct inference on the number of latent factors without making Gaussian assumptions. FGS provides a discussion of the (in)consistency of the PCA estimator with fixed $T$ from the vantage point of the well-known incidental parameter problem of the panel data literature (Neyman and Scott (1948); see Lancaster (2000) for a review). In PCA, sphericity allows to identify the number $k$ of factors from the first $k$ eigenvalue spacings being larger than zero, and being zero for the subsequent ones. If sphericity fails, all eigenvalue spacings are different from zero, and identification fails. On the contrary, the FA strategy does not exploit eigenvalue spacings and does not require sphericity. However, FA inference with small $T$ up to now mostly relies on (often restrictive) assumptions such as Gaussian variables (with a notable exception by Anderson and Amemiya (1988)) and error homoskedasticity across sample units. Those are untenable assumptions in our application to stock returns. The strong assumption of sphericity might also fail to hold in some samples, e.g., due to common ARCH effects in idiosyncratic errors (see Section 2). If it happens, our Monte Carlo experiments with non-Gaussian errors show that the eigenvalue spacing test of FGS exhibits size distortions of over 80 percentage points for a nominal size of 5%. That massive over-rejection translates into an average estimated number of factors often above 10 instead of 2 (true value) in our simulation design when $T=24$ (a sample size close to $T=20$ in our empirics). We get similar Monte Carlo results for a constrained Likelihood Ratio test when sphericity does not hold (see Section 5.3). Our Monte Carlo experiments also reveal that the chi-square test of the classical FA theory obtained under cross-sectionally homoskedastic Gaussian errors (Anderson (1963)) and implemented in most numerical packages suffers from massive over-rejection by around 80 percentage points for a nominal size of 5% when $T=24$.
@@
A central and practical issue in applied work with latent factors is to determine their number. For models with unobservable factors only, Connor and Korajczyk (1993) are the first to develop a test for the number of factors for large balanced panels of individual stock returns under covariance stationarity and homoskedasticity. Unobservable factors are estimated by the method of asymptotic principal components developed by Connor and Korajczyk (1986) (see also Stock and Watson (2002)). In heteroskedastic settings, a first strand of that literature focuses on consistent estimation (but not testing) procedures for the number of factors. Bai and Ng (2002) and Bai (2009) introduce a penalized least-squares strategy to estimate the number of factors. Ando and Bai (2015) extend that approach when explanatory variables are present in the linear specification. Onatski (2010) looks at the behavior of differences in adjacent eigenvalues to determine the number of factors when $n$ and $T$ are both large and comparable, while Ahn and Horenstein (2013) opt for a similar strategy based on eigenvalue ratios. Caner and Han (2014) propose an estimator with a group bridge penalization. Gagliardini, Ossola and Scaillet (2019) build a simple diagnostic criterion for approximate factor structure in large panels. Given observable factors, the criterion checks whether the errors are weakly cross-sectionally correlated, or share one or more unobservable common factors (interactive effects), and selects their number; see Gagliardini, Ossola and Scaillet (2020) for a survey of estimation of large dimensional conditional factor models in finance. A second strand of that literature develops inference procedures for hypotheses on the number of latent factors. Onatski (2009) deploys a characterization of the largest eigenvalues of a Wishart-distributed covariance matrix with large dimensions in terms of the Tracy-Widom Law. To get a Wishart distribution, Onatski (2009) assumes either Gaussian errors, or $T$ much larger than $n$. Under non-Gaussian errors or when $T$ is much lower than $n$, we also observe notable size distortions for such a test in our Monte Carlo simulations. The test described in Section 4 of Onatski (2009) yields overrejection often largely above 10 to 15 percentage points for a nominal size of 5% when $T=12,24$ for the data generating process used in Section 7. Kapetanios (2010) uses subsampling to estimate the limit distribution of the adjacent eigenvalues. Any of the above methods fail with $T$ small and non-Gaussian, non-spherical errors. In a fixed $T$ setting, Ahn et al. (2013) develop a test for the number of factors for a panel data model with random interactive effects and i.i.d. errors. The test statistic corresponds to the estimated value of the optimal Generalized Method of Moments (GMM) criterion (the usual $J$-test statistic of Hansen (1982)). Hayakawa et al. (2023) develop a likelihood ratio test in a transformed Gaussian quasi maximum likelihood approach for the number of factors for a dynamic panel data model with random interactive effects and i.i.d. spherical errors.
@@
This paper puts forward methodological and empirical contributions that complement the above literature. (i) On the methodological side, we extend the inferential tools of FA to non-Gaussian and non-i.i.d. settings. First, we characterize the asymptotic distribution of FA estimators obtained under a Pseudo Maximum Likelihood (PML) approach where the time-series dimension is held fixed while the cross-sectional dimension diverges. Hence, the asymptotic analysis targets short panels, and allows for cross-sectionally heteroskedastic and weakly dependent errors. Cochrane (2005, p. 226) argues in favour of the development of appropriate large-$n$ small-$T$ tools for evaluating asset pricing models, a problem only partially addressed in finance. In a short panel setting, Zaffaroni (2025) considers inference for latent factors in conditional linear asset pricing models under sphericity based on PCA, including estimation of the number of factors. Raponi, Robotti and Zaffaroni (2020) develop tests of beta-pricing models and a two-pass methodology to estimate the ex-post risk premia (Shanken (1992)) associated to observable factors (see Kleibergen and Zhan (2025) for robust-identification inference based a continuous updating GMM). Kim and Skoulakis (2018) deals with the error-in-variable problem of the two-pass methodology with small $T$ by regression-calibration under sphericity and a block-dependence structure. The small $T$ setting mitigates concerns for panel unbalancedness (outside the straightforward missing-at-random mechanism) and corresponds to a locally time-invariant factor structure accommodating globally time-dependent features of general forms. It is also appealing to macroeconomic data observed quarterly. For the sake of space, we put part of the theory, namely inference for FA estimates, in the Online Appendix (OA). We refer to Bai and Li (2016) for inference when $n$ and $T$ are both large (see Bai and Li (2012) for the cross-sectional independent case). Second, we build on our new theoretical results for FA to develop testing procedures for the number of latent factors in a short panel which rely on neither sphericity nor Gaussianity nor cross-sectional independence, thereby extending tests based on eigenvalues, as in Onatski (2009) under a spectral analysis targeting a generalized dynamic factor structure, to small $T$, and as in FGS, to non-spherical errors, thanks to an FA device. Such testing procedures have the practical advantage of not requiring a choice of penalization hyperparameters and of providing p-values to gauge statistical significance. Here, we deliver asymptotic feasible distributional theory, namely implementable theory based on estimated quantities instead of population (unknown) ones, even if FA residuals are not consistent estimates of the true errors under fixed $T$. We further derive the Asymptotically Uniformly Most Powerful Invariant (AUMPI) property of the FA likelihood ratio (LR) test statistic in the non-Gaussian case under inequality restrictions on the DGP parameters, and cover inference with weak factors. The AUMPI property is rare and sought-after in testing procedures (see Engle (1984) for a discussion and Romano, Shaikh and Wolf (2010) for a survey of optimality approaches in testing problems), and often holds only under restrictive assumptions such as Gaussianity. For the user of the test, it provides a theoretical guarantee of a statistical procedure with better power properties w.r.t. another choice of testing. We show that the AUMPI property can hold even if the asymptotic distribution of the LR statistic is driven by a weighted sum of independent chi-square variates in our context instead of a single chi-square variate. We achieve that by providing novel sufficient conditions so that the ratio of the density under local alternative hypotheses and the null hypothesis satisfies the Monotone Likelihood Ratio (MLR) property. Hence, the FA theory gathered in the below delivers a body of new results including their proofs, and differs completely from PCA theory. (ii) On the empirical side, we apply our FA methodology to panels of monthly U.S. stock returns with large cross-sectional and small time-series dimensions, and investigate how the number of driving factors changes over time and particular periods. Furthermore, month after month, we provide a novel separation of the risk coming from the systematic part and the risk coming from the idiosyncratic part of returns in short subperiods of bear vs. bull market based on the selected number of factors. Such a decomposition with PCA estimates is invalid without sphericity because of inconsistency of factor estimates. We observe an uptrend in the estimated paths of total and idiosyncratic volatilities (see also Campbell et al. (2023)) while the systematic risk explains a large part of the cross-sectional total variance in bear markets but is not driven by a single latent factor. We also investigate whether standard observed factors span the estimated latent factors using rank tests that suit our fixed $T$ setting. Observed factors, even beyond the traditional ones, struggle spanning latent factors with a discrepancy between the dimensions of the two factor spaces decreasing over time.
@@
The outline of the paper is as follows. In Section 2, we consider a linear latent factor model and introduce test statistics on the number of latent factors based on FA. Section 3 lists the regularity assumptions underlying the theoretical results. Section 4 presents a feasible asymptotic distributional theory for inference in short panels under a block-dependence structure to allow for weak dependence in the cross-section. Sections 5 and 6 develop further theory supporting the advantages of our method. Section 5 discusses three special cases, i.e., Gaussian errors (yielding the classical chi-square test of FA theory in the cross-sectionally homoskedastic case), non-Gaussian settings where the same asymptotic distribution as under Gaussian errors still holds for the test statistics, and spherical errors (leading to PCA). It considers the impact for test validity of wrongly assuming Gaussian errors or sphericity. Section 6 is dedicated to local asymptotic power and AUMPI properties. Section 7 gives a Monte Carlo assessment of size and power and selection procedure for the number of factors for the LR test. We provide our empirical application in Section 8. We put our concluding remarks in Section 9. In the Online Appendix (OA), we gather proofs of the main theoretical results in Appendix A as well as proofs of Lemmas supporting them. We place additional theory and numerical checks in Appendices B and C. Appendix D collects the maximum value of $k$ as a function of $T$. We gather all explicit formulas not listed in the core text but useful for coding in an online “Supplementary Materials for Coding” (SMC) attached to the replication files. We also put there other numerical checks and additional Monte Carlo results to assess the impact of non-sphericity for the eigenvalue spacing test of FGS and the constrained LR test, as well as of using the classical chi-square test on the size and selection procedure for the number of factors.
@@ 2 Latent factor model and likelihood ratio test
where $y_{i}=(y_{i,1},...,y_{i,T})^{\prime}$ and $\varepsilon_{i}=(\varepsilon_{i,1},...,\varepsilon_{i,T})^{\prime}$ are $T$-dimensional vectors of observed data and unobserved error terms for individual $i$. The $k$-dimensional vectors $\beta_{i}=(\beta_{i,1},...,\beta_{i,k})^{\prime}$ are latent individual effects (non-random incidental parameters), while $\mu$ and $F$ are a $T\times 1$ vector and a $T\times k$ matrix of unknown parameters. The number of latent factors $k$ is an unknown integer smaller than $T$. In matrix notation, model (1) reads $Y=\mu 1_{n}^{\prime}+F\beta^{\prime}+\varepsilon$, where $Y$ and $\varepsilon$ are $T\times n$ matrices, $\beta$ is the $n\times k$ matrix with rows $\beta_{i}^{\prime}$, and $1_{n}$ is a $n$-dimensional vector of ones.
@@
Matrix $V_{\varepsilon}$ is the limit cross-sectional average of the unconditional covariance matrix of the - possibly heterogeneous - errors. The diagonality condition in Assumption 1 is standard in FA (in the more restrictive formulation involving i.i.d. data). Assumption 1 allows for serial dependence in idiosyncratic errors in the form of martingale difference sequences, like individual GARCH and Stochastic Volatility (SV) processes, as well as weak cross-sectional dependence (see Assumption 2 below). It also accommodates common time-varying components in idiosyncratic volatilities by allowing different entries along the diagonal of $V_{\varepsilon}$; see Renault, Van Der Heijden and Werker (2023) for arbitrage pricing in such settings. When there is a common random component in idiosyncratic volatilities, i.e., an idiosyncratic volatility factor, we have $V_{\varepsilon}=\underset{n\rightarrow\infty}{plim}\frac{1}{n}\varepsilon\varepsilon^{\prime}$ by a suitable version of the Law of Large Number (LLN) conditional on the sigma-field generated by this common component. With fixed $T$, we treat the sample realizations of the common component in idiosyncratic volatilities as unknown time fixed effects (the diagonal elements of matrix $V_{\varepsilon}$), which yields time heterogeneous distributions for the errors. It is how the unconditional expectation in Assumption 1 has to be understood. The assumption of zero off-diagonal elements in matrix $V_{\varepsilon}$ rules out pervasive serial autocorrelation in the errors, i.e., $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}E[\varepsilon_{i,t}\varepsilon_{i,s}]=0$ for $t\neq s$. Generalizations to non-diagonal $V_{\varepsilon}$ are possible (Jöreskog (1970)), but stronger restrictions are required to maintain nonnegative degrees of freedom (see the discussion following (FA1) and (FA2) below). Besides, numerical optimisation becomes less tractable. Such an extension might be needed if we want to deal with serially correlated errors possibly present in macroeconomic panels and we can modify the theory developed in this paper to suit such a setting.
In our empirics with a large cross-sectional panel of returns for $n$ assets over a short time span with $T$ periods, vectors $y_{i}$ and $\varepsilon_{i}$ stack the monthly returns and the idiosyncratic errors of stock $i$. Any row vector $f_{t}^{\prime}:=(f_{t,1},...,f_{t,k})$ of matrix $F$ yields the latent factor values in a given month $t$, and vector $\beta_{i}$ collects the factor loadings of stock $i$. In this finance application, we assume the No-Arbitrage (NA) principle to hold, so that the entries $\mu_{t}$ in the intercept vector in Equation (1) account for the (possibly time-varying) risk-free rate and (possibly non-zero) cross-sectional mean of stock betas. Specifically, under NA and assuming that the assets are embedded into a continuous economy, the intercept term in the asset return model $y_{i}=\mu_{i}+\tilde{F}\tilde{\beta}_{i}+\varepsilon_{i}$ is $\mu_{i}=r_{f}+1_{T}\nu^{\prime}\tilde{\beta}_{i}$, where $r_{f}$ is the $T$-dimensional vector whose entries collect the (possibly time-varying) risk-free rates, $\nu=(\nu_{1},...,\nu_{k})^{\prime}$ is a $k$-dimensional vector of parameters, and $1_{T}$ is a $T$-dimensional vector of ones (see e.g. the derivation in Gagliardini, Ossola and Scaillet (2016) for an Arbitrage Pricing Theory (Ross (1976)) under weak cross-sectional dependence (Chamberlain and Rothschild (1983))). We can absorb term $1_{T}\nu^{\prime}\tilde{\beta}_{i}$ into the systematic part to get $y_{i}=r_{f}+F\tilde{\beta}_{i}+\varepsilon_{i}$ with $F=\tilde{F}+1_{T}\nu^{\prime}$. It holds irrespective of the latent factors being tradable or not. If the factors are tradable, we further have $\nu=0$ from the NA restriction. Akin to standard formulation of FA, we recenter the latent effects by subtracting their mean $\tilde{\mu}_{\tilde{\beta}}=\frac{1}{n}\sum_{i=1}^{n}\tilde{\beta}_{i}$, to get model (1) with $\beta_{i}=\tilde{\beta}_{i}-\tilde{\mu}_{\tilde{\beta}}$ and $\mu=r_{f}+F\tilde{\mu}_{\tilde{\beta}}$. Instead, Chamberlain and Rothschild (1983) derive the APT with countably many assets and show that pricing errors $\zeta_{i}$ are square summable under NA. Then, we get $y_{i}=\mu+1_{T}\zeta_{i}+F\beta_{i}+\varepsilon_{i}$ with $\sum_{i=1}^{n}\zeta_{i}^{2}\leq C$, for all $n$ and a constant $C$. Our distributional results on FA estimators and LR test statistic are unaffected by the presence of $\zeta_{i}$, because we can show that the estimated $T\times T$ covariance matrix of excess returns is unchanged up to terms of order $o_{p}(n^{-1/2})$. If the NA principle is violated, but the elements of the vector $\mu_{i}$ are constant, i.e., $\mu_{i}=\alpha_{i}1_{T}$ for some scalar $\alpha_{i}$, then the vector $1_{T}$ becomes an additional factor in the model. In any case, our starting point is the linear FA model in (1): $y_{i,t}=\mu_{t}+f_{t}^{\prime}\beta_{i}+\varepsilon_{i,t}$, that is the standard formulation in asset pricing. We cover the Capital Asset Pricing Model (CAPM) when the single latent factor is the excess return of the market portfolio.